Marine, Inc., manufactures a product that is available in both a flexible and a rigid model. The company has made the rigid mcame back with the recommendation that four activity cost pools be used. These cost pools and their associated activities areReq 1A Reg 1B Reg 2 Req 3A Req 3B Compute the unit product cost for one unit of each model. Unit Product Cost Flexible Rigidis question by entering your answers in the tabs below. Req 1A Req 1B Req 2 Req ЗА Req 3B Compute the activity rate for eachReq 1A Req IA Req 18 Reg 1B Reg 2 Req 2 Req ЗА Reg 3A Req 3B Req 3B Using activity-based costing, determine the total amountReq 1A Req 1B Reg 2 Req 3A Req 3B Using activity-based costing, compute the unit product cost for one unit of each model. (Do
Answer #1

Solution 1:

Direct labor hours estimated = 2000*3 + 9000*1.50 = 19500 hours

Predetermined overhead rate = $741000 / 19500 = $38 per DLH

Solution 2:

Computation of Unit Product Cost
Flexible Rigid
Direct Materials $100.00 $75.00
Direct Labor $50.00 $35.00
Overhead Cost per unit ($38*Labor hours per unit) $114.00 $57.00
Unit product cost $264.00 $167.00

Solution 3:

Computation of Activity rates
Activity cost pool Overhead costs Estimated activity Activity Rate
Purchase orders $26,500 530 $50.00 per order
Rework requests $16,500 330 $50.00 per request
Product testing $3,00,000 2400 $125.00 per test
Machine Related $3,98,000 3980 $100.00 per MH

Solution 4:

Computation of Overhead allocated and Overhead per unit
Flexible Rigid
Activity cost pool Activity Rate Estimated Activity Overhead Allocated Estimated Activity Overhead Allocated
Purcahse orders $50.00 140 $7,000 390 $19,500
Rework requests $50.00 55 $2,750 275 $13,750
Product testing $125.00 650 $81,250 1750 $2,18,750
Machine Related $100.00 1100 $1,10,000 2880 $2,88,000
Total Overhead allocated $2,01,000 $5,40,000
No. of units 2000 9000
Overhead per unit 100.50 60.00

Solution 5:

Computation of Unit Product Cost
Flexible Rigid
Direct Materials $100.00 $75.00
Direct Labor $50.00 $35.00
Overhead Cost per unit $100.50 $60.00
Unit product cost $250.50 $170.00